### v2

@@ 1 Introduction
Existing Deepfake detection methods primarily exploit specific forgery traces within forged face data, such as blurred edges Shiohara & Yamasaki (2022), spatial inconsistencies Zhang et al. (2025a), and temporal jitter Xu et al. (2025a), to obtain a discriminative detection result. With the rapid advancement of generative models, particularly Large Language Models (LLMs) and Vision-Language Models (VLMs), new identification demands have emerged for Deepfake detection and reasoning. Several recent methods have attempted to leverage LLMs for explainable Deepfake detection. For example, M2F2-Det Guo et al. (2025) enhances the explainability of Vicuna-7B Chiang et al. (2023) by integrating the forgery prompt learning and the CLIP’s multimodal representation ability. KFD Yu et al. (2025) proposes aligning face images with textual descriptions to generate forgery consistency maps, which achieves multimodal alignment through prompt learning. RAIDX Li et al. (2025a) integrates retrieval-augmented generation and group relative policy optimization to enhance the detection accuracy and decision explainability of LLMs. The aforementioned methods primarily rely on combining prompt engineering with LLMs, which enhance the ability of explainable Deepfake detection through elaborate prompt design. In this paper, we focus on directly leveraging VLMs for Deepfake detection. By integrating reinforcement learning and multimodal alignment of reasoning content, we aim to enhance the accuracy and reliability of VLMs in Deepfake detection and reasoning. VLMs exhibit superior performance in associating textual and visual modalities, thus generating more reliable content by maintaining multimodal representation alignment during the reasoning process Zhai et al. (2024). While existing pre-trained VLMs perform well in scenarios involving strong semantic signals, such as Visual Question Answering and Image Captioning, they tend to underperform in the task of Deepfake detection scenario. The primary reason is that forged face images are often visually difficult to distinguish from genuine ones, since the forgery traces are typically subtle or imperceptible. Therefore, it is necessary to enhance existing pre-trained VLMs for the specific scenario of Deepfake detection.
@@ 2 Related Works
Deepfake detection. Deepfake detection primarily focuses on binary classification, which aims to discriminate the authenticity of a given image or video. Various effective strategies have been proposed, including data augmentation Yin et al. (2024), biometric information analysis Xu et al. (2023), and attention mechanisms Zhao et al. (2021). Besides, some methods have also investigated the spatial localization task and temporal localization task to cope with different detection scenarios Zhang et al. (2025a); Xu et al. (2025a). These methods basically provide a discriminative decision for a given face data, lacking additional forensic information for humans. Recently, researchers have explored utilizing LLMs to generate the textual reasoning content about forgery traces, which enables the Deepfake detection to be more explainable Guo et al. (2025); Yu et al. (2025). The LLM-based methods typically fine-tune models using image-text pairs, where the answers usually consist of only text information about forgery traces. It should be noted that a conventional Deepfake detector is often employed as an expert network to guide the LLM’s reasoning process. In this paper, MARE introduces additional spatial information to motivate the VLM to generate more reliable reasoning content with text-spatial alignment.
@@
Figure 2: Diagrammatic overview of the proposed MARE framework. The reward functions and RLHF algorithm are displayed on the right side (see Section 3.3 for details). The vision encoder, text tokenizer, and large language model decoder are key components of VLM. The forgery disentanglement module is introduced for forgery traces extraction(see Section 3.4 for details). During the inference phase, MARE generates text-spatially aligned reasoning content for a given image-text query.
@@
Multimodal alignment in VLMs. Alignment is crucial for improving the downstream task accuracy, safety, and reliability of VLMs Li et al. (2025b). The general alignment algorithm is reinforcement learning from human feedback (RLHF). RLHF is a machine learning technique that uses human feedback to train and align models for performing reasoning tasks more aligned with human preferences Li et al. (2024); Xie et al. (2025). The core principle of RLHF is to construct human feedback annotation data and design reward mechanisms. MM-RLHF Zhang et al. (2025b) constructs a high-quality, fine-grained dataset with 120k preference comparison pairs to advance the alignment of VLMs. Group relative policy optimization utilizes rule-based reward to improve the models’ chain-of-thought ability for more accurate and reliable reasoning content Shao et al. (2024). RLHF has become one of the most popular and effective paradigms to align VLMs. However, existing LLM-based Deepfake detectors have insufficient exploration of alignment Meng et al. (2025). In this paper, MARE initially constructs a preference text-spatially aligned dataset. Based on these data, MARE introduces elaborate reward functions to improve the accuracy and reliability of VLMs in Deepfake decision attribution.
@@ 3.1 Overview
To enhance the accuracy and reliability of VLMs in Deepfake detection and reasoning, we propose MARE under the RLHF paradigm. The diagrammatic overview is shown in Figure 2. We design comprehensive reward functions and construct a human-preference dataset to incentivize the VLMs to generate text-spatially aligned reasoning content. Meanwhile, MARE introduces a forgery disentanglement module (FDM) to capture intrinsic forgery traces for precise authenticity detection.
@@
In the inference phase, MARE generates text-spatially aligned reasoning content $\langle\hat{t}_{i},\hat{b}_{i}\rangle$ for a given query $\langle v_{i},q_{i}\rangle$, where $\hat{t}_{i}$ is the textual reasoning content that contains the authenticity detection result and related explanation, $\hat{b}_{i}$ provides the spatial localization of regions mentioned in $\hat{t}_{i}$. $\hat{t}_{i}$ and $\hat{b}_{i}$ provide mutual validation to improve the accuracy and reliability of reasoning content.
@@ 3.2 Multimodal Alignment
Alignment could improve the accuracy, safety, and reliability of VLMs in the downstream task Li et al. (2025b). In the Deepfake forensic domain, it is particularly crucial to enhance the accuracy and reliability of reasoning content through multimodal alignment. In this paper, the objective of multimodal alignment is to incentivize the VLM to generate reasoning content that aligns textual and spatial information. For instance, if the text describes forgery traces in the eyes, the reasoning output should provide the corresponding bounding boxes for the eyes region. This mechanism could not only mitigate the VLM’s hallucinations but also enhance the reliability of the reasoning content.
We construct the Deepfake multimodal alignment dataset (DMA) by augmenting the existing image-text dataset Zhang et al. (2024); Yu et al. (2025), denoted as $\mathcal{D}_{src}=\left\{\langle v_{i},q_{i},t_{i}\rangle\right\}_{i=1}^{N}$, where $v_{i}$ indicates the face image, $q_{i}$ is a query question, and $t_{i}$ is the ground-truth textual description about authenticity and detailed analysis. We propose a two-stage extraction and localization pipeline. Specifically, a set of facial regions $\mathcal{P}_{key}$ is defined (e.g., eyes, nose, or mouth, see appendix B for details). We perform keyword retrieval $\mathbb{E}(\cdot,\cdot)$ for each $t_{i}$ to extract the specific regions $r_{i}$. Then, a face landmark detection model $\mathbb{L}(\cdot,\cdot)$ is utilized to locate the spatial bounding boxes $b_{i}$ from the face image $v_{i}$. The pipeline is formulated as:
@@ 3.3 Reinforcement Learning Reward Design
Format reward. Format reward is to check whether the VLM’s responses follow the human preferences format. Specifically, MARE requires the model to present the reasoning paths enclosed within <think>...</think> tags. The final reasoning result within <answer>...</answer> tags should contain two parts: a "explanation" field with a string value describing why the face image is real or fake, and a "bboxes" field listing the bounding box of each region mentioned in the "explanation" that serves as evidence. Formally,
@@
Accuracy reward. To support the Deepfake detection and reasoning task, MARE introduces an accuracy reward $\mathcal{R}_{a}$ based on the correctness of authenticity prediction within the response. Specifically, we extract keyword "fake" or "real" from the "explanation" field. Serving as a basic and essential signal, $\mathcal{R}_{a}$ is to incentivize VLMs explore a valid chain-of-thought path for precise Deepfake detection. Formally,
@@
Alignment reward. The alignment reward $\mathcal{R}_{align}$ is designed to incentivize the VLM to generate text-spatially aligned reasoning content. This mechanism helps to mitigate the VLM’s hallucinations. Then MARE could generate more reliable Deepfake detection and reasoning results. Specifically, for a candidate response $o_{k}$, we extract the set of facial regions $\mathcal{P}_{\hat{b}_{i}}$ from "bboxes" field, and the set of mentioned facial regions $\mathcal{P}_{\hat{t}_{i}}$ from the "explanation" field. The higher the overlap between these two sets, the greater the alignment between textual and spatial reasoning content. Formally,
$\mathcal{R}_{\text{align }}\left(o_{k}\right)=\frac{\left|\mathcal{P}_{\hat{b}_{i}}\cap\mathcal{P}_{\hat{t}_{i}}\right|}{\left|\mathcal{P}_{\hat{b}_{i}}\cup\mathcal{P}_{\hat{t}_{i}}\right|+\epsilon}$ (6)
@@ 3.4 Forgery Disentanglement Module
Specifically, FDM first utilizes a backbone to obtain the shared latent representations $\mathcal{F}$ of the given face image. $\mathcal{F}$ are disentangled into identity features $\mathcal{F}_{I}$, structural features $\mathcal{F}_{s}$, and forgery trace features $\mathcal{F}_{f}$. To facilitate feature disentanglement, the FDM incorporates dedicated classifiers and loss functions to guide FDM training. The identity classifier utilizes $\mathcal{F}_{I}$ to generate identity predictions, then calculates the identity prediction loss $\mathcal{L}_{I}$ based on the ground truth identity annotations. Focal loss is utilized to guide the model to focus on challenging samples. Formally,
@@
Besides identity information and forgery traces, the face image also contains structural information, such as illumination, expressions, and posture information. To prevent these factors from interfering with forgery detection, they are disentangled into structural features $\mathcal{F}_{s}$. Furthermore, $\mathcal{F}_{s}$, $\mathcal{F}_{I}$, and $\mathcal{F}_{f}$ are concatenated for initial feature reconstruction. This imposes an information constraint that maintains the stability of the disentanglement framework. Formally,
$\mathcal{L}_{r}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{F}_{i}-\mathcal{F}_{i}^{(r)}\right)^{2}$ (9)
where $\mathcal{F}_{i}$ denotes the shared latent representations within Figure 4, and $\mathcal{F}_{i}^{(r)}$ is the reconstruction features.
@@ 3.5 Training and Inference
Training. MARE contains two training stages: supervised training for the FDM and reinforcement learning-based optimization for the VLM. First, the FDM is trained based on existing Deepfake datasets. The overall loss function is defined as follows:
@@
where $\lambda_{*}$ is the weight parameter to balance each loss component. Secondly, MARE fine-tunes the VLM based on the constructed DMA dataset via a reinforcement learning algorithm. The primary objective is to incentivize the VLM’s reasoning capabilities in Deepfake detection and reasoning. The final reward function $\mathcal{R}$ is defined as follows:
@@
where $\beta_{*}$ denotes the weight parameter.
@@ 4.3 Reasoning performance
Quantitative result. To quantitatively evaluate the generated textual description, we extract identification keywords (e.g., "fake" or "real") from the reasoning content. Then, Acc and F1 are calculated based on the ground truth labels. The identification performance of MARE and comparison methods is shown in Table 3.
The pre-trained Qwen Bai et al. (2025), InternVL Chen et al. (2024), and LLaVA Liu et al. (2023) are utilized as baseline models. The results indicate the pre-trained VLMs exhibit suboptimal performance on Deepfake detection and reasoning. The primary reason is that forged face images are often visually difficult to distinguish from genuine ones, since the forgery traces are typically imperceptible. After supervised fine-tuning (SFT) with annotated data, LLaVA achieves significant improvements in both Acc and F1 scores. DDVQA-BLIP Zhang et al. (2024) and M2F2 combined with SFT to enhance the reasoning capability of VLMs, which achieves further performance improvements. However, the alignment problem in the reasoning process has not been sufficiently explored. MARE constructs a multimodal alignment dataset and designs comprehensive reward functions, and utilizes reinforcement learning to incentivize the VLM to generate accurate and reliable reasoning content. That is the reason that MARE achieves superior performance on both Acc and F1 scores.
@@
image look fake/real?". Qwen identifies all samples as "real", and generates the textual description that has a generalized and templated style. GPT-5 classifies all samples as "fake". Compared to Qwen, the textual description generated by GPT-5 captures more details, such as eyes, nose, and teeth edges. The results in Figure 5 and Table 3 indicate that the existing general-purpose VLMs have unsatisfactory accuracy in face image authenticity detection. M2F2 achieves improved detection accuracy by introducing expert knowledge. Besides, M2F2 generates fine-grained textual descriptions based on a detailed prompt (e.g., Figure 5 (a) utilizes "The mouth in the image is fake, tell me why..." as the prompt). In contrast, MARE produces reasoning content with text-spatial alignment, which is more aligned with human preferences. It should be noted that MARE uses the prompt "Does this image look fake/real...?", which does not mention specific facial regions. Despite this, MARE could still perceive, locate, and analyze the detected face image, thus generating more accurate and reliable reasoning content.

### v3

@@ 1 Introduction
Existing Deepfake detection methods primarily exploit specific forgery traces within forged face data, such as blurred edges Shiohara and Yamasaki (2022), spatial inconsistencies Zhang et al. (2025a), and temporal jitter Xu et al. (2025a), to obtain a discriminative detection result. With the rapid advancement of generative models, particularly Large Language Models (LLMs) and Vision-Language Models (VLMs), new identification demands have emerged for Deepfake detection and reasoning. Several recent methods have attempted to leverage LLMs for explainable Deepfake detection. For example, M2F2-Det Guo et al. (2025) enhances the explainability of Vicuna-7B Chiang et al. (2023) by integrating the forgery prompt learning and the CLIP’s multimodal representation ability. KFD Yu et al. (2025) proposes aligning face images with textual descriptions to generate forgery consistency maps, which achieves multimodal alignment through prompt learning. RAIDX Li et al. (2025a) integrates retrieval-augmented generation and group relative policy optimization to enhance the detection accuracy and decision explainability of LLMs. The aforementioned methods primarily rely on combining prompt engineering with LLMs, which enhance the ability of explainable Deepfake detection through elaborate prompt design. In this paper, we focus on directly leveraging VLMs for Deepfake detection. By integrating reinforcement learning and multimodal alignment of reasoning content, we aim to enhance the accuracy and reliability of VLMs in Deepfake detection and reasoning. VLMs exhibit superior performance in associating textual and visual modalities, thus generating more reliable content by maintaining multimodal representation alignment during the reasoning process Zhai et al. (2024). While existing pre-trained VLMs perform well in scenarios involving strong semantic signals, such as Visual Question Answering and Image Captioning, they tend to underperform in the task of Deepfake detection scenario. The primary reason is that forged face images are often visually difficult to distinguish from genuine ones, since the forgery traces are typically subtle or imperceptible. Therefore, it is necessary to enhance existing pre-trained VLMs for the specific scenario of Deepfake detection.
@@ 2 Related Work
Deepfake detection. Deepfake detection primarily focuses on binary classification, which aims to discriminate the authenticity of a given image or video Yan et al. (2025). Various effective strategies have been proposed, including data augmentation Yin et al. (2024), biometric information analysis Xu et al. (2023), and attention mechanisms Zhao et al. (2021). Besides, some methods have also investigated the spatial localization task and temporal localization task to cope with different detection scenarios Zhang et al. (2025a); Xu et al. (2025a). These methods basically provide a discriminative decision for a given face data, lacking additional forensic information for humans. Recently, researchers have explored utilizing LLMs to generate the textual reasoning content about forgery traces, which enables the Deepfake detection to be more explainable Guo et al. (2025); Yu et al. (2025). The LLM-based methods typically fine-tune models using image-text pairs, where the answers usually consist of only text information about forgery traces. It should be noted that a conventional Deepfake detector is often employed as an expert network to guide the LLM’s reasoning process. In this paper, MARE introduces additional spatial information to motivate the VLM to generate more reliable reasoning content with text-spatial alignment.
@@
Figure 2: Diagrammatic overview of the proposed MARE framework. The reward functions and RLHF algorithm are displayed on the right side (see Section 3.3 for details). The vision encoder, text tokenizer, and large language model decoder are three key components of VLM. The forgery disentanglement module is introduced for intrinsic forgery traces extraction(see Section 3.4 for details). During the inference phase, MARE generates text-spatially aligned reasoning content for a given image-text query.
@@
Multimodal alignment in VLMs. Alignment is crucial for improving the downstream task accuracy, safety, and reliability of VLMs Li et al. (2025b). The general alignment algorithm is reinforcement learning from human feedback (RLHF). RLHF is a machine learning technique that uses human feedback to train and align models for performing reasoning tasks more aligned with human preferences Li et al. (2024); Xie et al. (2025). The core principle of RLHF is to construct human feedback annotation data and design reward mechanisms. MM-RLHF Zhang et al. (2025b) constructs a high-quality, fine-grained dataset with 120k preference comparison pairs to advance the alignment of VLMs. Group relative policy optimization utilizes rule-based reward to improve the models’ chain-of-thought ability for more accurate and reliable reasoning content Shao et al. (2024). RLHF has become one of the most popular and effective paradigms to align VLMs. However, existing LLM-based Deepfake detectors have insufficient exploration of alignment Meng et al. (2025). In this paper, MARE initially constructs a preference text-spatially aligned dataset. Based on these data, MARE introduces elaborate reward functions to improve the accuracy and reliability of VLMs in Deepfake detection and reasoning.
@@ 3.1 Overview
To enhance the accuracy and reliability of VLMs in Deepfake detection and reasoning, we propose MARE under the RLHF paradigm. The diagrammatic overview is shown in Figure 2. We design comprehensive reward functions and construct a human-preference dataset to incentivize the VLMs to generate text-spatially aligned reasoning content. Meanwhile, MARE introduces a forgery disentanglement module (FDM) to capture intrinsic forgery traces for precise authenticity identification.
@@
In the inference phase, MARE generates text-spatially aligned reasoning content $\langle\hat{t}_{i},\hat{b}_{i}\rangle$ for a given query $\langle v_{i},q_{i}\rangle$, where $\hat{t}_{i}$ is the textual reasoning content that contains the authenticity detection result and related explanation, $\hat{b}_{i}$ provides the spatial bounding boxes of regions mentioned in $\hat{t}_{i}$. $\hat{t}_{i}$ and $\hat{b}_{i}$ provide mutual validation to improve the accuracy and reliability of reasoning content.
@@ 3.2 Multimodal Alignment
Alignment could improve the accuracy, safety, and reliability of VLMs in the downstream task Li et al. (2025b). In the Deepfake forensic domain, it is particularly crucial to enhance the accuracy and reliability of reasoning content through multimodal alignment. In this paper, the objective of multimodal alignment is to incentivize the VLM to generate reasoning content that aligns textual and spatial information. For instance, if the text describes forgery traces about mouth, the reasoning output should provide corresponding bounding boxes for the mouth region. This mechanism could not only mitigate the VLM’s hallucinations but also enhance the reliability of the reasoning content.
Specifically, we construct the Deepfake multimodal alignment dataset (DMA) by augmenting the existing image-text dataset Zhang et al. (2024); Yu et al. (2025), denoted as $\mathcal{D}_{src}=\left\{\langle v_{i},q_{i},t_{i}\rangle\right\}_{i=1}^{N}$, where $v_{i}$ indicates the face image, $q_{i}$ is a query question, and $t_{i}$ is the ground-truth textual description about authenticity and detailed analysis. We propose a two-stage extraction and localization pipeline. Specifically, a set of facial regions $\mathcal{P}_{key}$ is defined (e.g., eyes, nose, or mouth, see appendix B for details). We perform keyword retrieval $\mathbb{E}(\cdot,\cdot)$ for each $t_{i}$ to extract the specific regions $r_{i}$. Then, a face landmark detection model $\mathbb{L}(\cdot,\cdot)$ is utilized to locate the spatial bounding boxes $b_{i}$ from the face image $v_{i}$. The pipeline is formulated as:
@@ 3.3 Reinforcement Learning Reward Design
Format reward. Format reward is to check whether the VLM’s responses follow the human preferences format. Specifically, MARE requires the model to present the reasoning paths enclosed within <think>...</think> tags. The final reasoning result within <answer>...</answer> tags should contain two parts: an "explanation" field with a string value describing why the face image is real or fake, and a "bboxes" field listing the bounding box of each region mentioned in the "explanation" that serves as evidence. Formally,
@@
Accuracy reward. To support the Deepfake detection and reasoning task, MARE introduces an accuracy reward $\mathcal{R}_{a}$ based on the correctness of authenticity prediction within the response. Specifically, we extract keyword "fake" or "real" from the "explanation" field. Serving as a basic and essential signal, $\mathcal{R}_{a}$ is to incentivize VLMs explore a valid chain-of-thought path for accurate Deepfake detection. Formally,
@@
Alignment reward. The alignment reward $\mathcal{R}_{align}$ is designed to incentivize the VLM to generate text-spatially aligned reasoning content. Then MARE could generate more reliable Deepfake detection and reasoning results. Specifically, for a candidate response $o_{k}$ of $\mathcal{D}_{ma}^{(i)}$, we extract the set of facial regions $\mathcal{P}_{\hat{b}_{i}^{(k)}}$ from "bboxes" field, and the set of mentioned facial regions $\mathcal{P}_{\hat{t}_{i}^{(k)}}$ from the "explanation" field. The higher the overlap between these two sets, the greater the alignment between textual and spatial reasoning content. Formally,
$\mathcal{R}_{align}\left(o_{k}\right)=\frac{\left|\mathcal{P}_{\hat{b}_{i}^{(k)}}\cap\mathcal{P}_{\hat{t}_{i}^{(k)}}\right|}{\left|\mathcal{P}_{\hat{b}_{i}^{(k)}}\cup\mathcal{P}_{\hat{t}_{i}^{(k)}}\right|+\epsilon}$ (6)
@@ 3.4 Forgery Disentanglement Module
Specifically, FDM first utilizes a backbone to obtain the shared latent representations $\mathcal{F}^{(i)}$ of the given face image $v_{i}$. $\mathcal{F}^{(i)}$ are disentangled into identity features $\mathcal{F}_{I}$, structural features $\mathcal{F}_{s}$, and forgery trace features $\mathcal{F}_{f}$. To facilitate feature disentanglement, the FDM incorporates dedicated classifiers and loss functions to guide FDM training. The identity classifier utilizes $\mathcal{F}_{I}$ to generate identity predictions, then calculates the identity prediction loss $\mathcal{L}_{I}$ based on the ground truth identity annotations. Focal loss is utilized to guide the model to focus on challenging samples. Formally,
@@
Besides identity information and forgery traces, the face image also contains structural information, such as illumination, expressions, and head posture information. To prevent these factors from interfering with forgery detection, they are disentangled into structural features $\mathcal{F}_{s}$. Furthermore, $\mathcal{F}_{s}$, $\mathcal{F}_{I}$, and $\mathcal{F}_{f}$ are concatenated for initial feature reconstruction. The reconstruction loss $\mathcal{L}_{r}$ imposes an information constraint that maintains the stability of the disentanglement framework. Formally,
$\mathcal{L}_{r}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{F}^{(i)}-\hat{\mathcal{F}}^{(i)}\right)^{2}$ (9)
where $\mathcal{F}^{(i)}$ denotes the shared latent representations within Figure 4, and $\hat{\mathcal{F}}^{(i)}$ is the reconstruction features.
@@ 3.5 Training and Inference
Training. MARE contains two training stages: supervised training for the FDM and RL-based optimization for the VLM. First, the FDM is trained based on existing Deepfake datasets. The overall loss function is defined as follows:
@@
where $\lambda_{1}=\lambda_{3}=1\times 10^{-4}$ and $\lambda_{2}=1$ are the weight parameter to balance each loss component. Secondly, MARE fine-tunes the VLM based on the constructed DMA dataset via a reinforcement learning algorithm. The primary objective is to incentivize the VLM’s reasoning capabilities in Deepfake detection and reasoning. The final reward function $\mathcal{R}$ is defined as follows:
@@
where $\beta_{2}=0.6$ and $\beta_{1}=\beta_{3}=\beta_{4}=\beta_{5}=0.1$ denote the weight parameter.
@@ 4.3 Reasoning Performance
Quantitative performance. To quantitatively evaluate the generated textual description, we extract identification keywords (e.g., "fake" or "real") from the reasoning content. Then, Acc and F1 are calculated based on the ground truth labels. The identification performance of MARE and comparison methods is shown in Table 3.
The pre-trained Qwen Bai et al. (2025), InternVL Chen et al. (2024), and LLaVA Liu et al. (2023) are utilized as baseline models. The results indicate the pre-trained VLMs underperform on Deepfake detection and reasoning. The primary reason is that forged face images are often visually difficult to distinguish from genuine ones, since the forgery traces are typically imperceptible. After supervised fine-tuning (SFT) with annotated data, LLaVA achieves significant improvements in both Acc and F1 scores. DDVQA-BLIP Zhang et al. (2024) and M2F2 combined with SFT to enhance the reasoning capability of VLMs, which achieves further performance improvements. However, the alignment problem in the reasoning process has not been sufficiently explored. MARE constructs a multimodal alignment dataset and designs comprehensive reward functions, and utilizes reinforcement learning to incentivize the VLM to generate accurate and reliable reasoning content. That is the reason that MARE achieves superior performance on both Acc and F1 scores.
@@
image look fake/real?". Qwen identifies all samples as "real", and generates the textual description that has a generalized and templated style. GPT-5 classifies all samples as "fake". Compared to Qwen, the textual description generated by GPT-5 captures more details, such as eyes, nose, and teeth edges. The results in Figure 5 and Table 3 indicate that the existing general-purpose VLMs have unsatisfactory accuracy in face image authenticity identification. M2F2 achieves improved detection accuracy by introducing expert knowledge. Besides, M2F2 generates fine-grained textual descriptions based on a detailed prompt (e.g., Figure 5 (a) utilizes "The mouth in the image is fake, tell me why..." as the prompt). In contrast, MARE produces reasoning content with text-spatial alignment, which is more aligned with human preferences. It should be noted that MARE uses the prompt "Does this image look fake/real...?", which does not mention specific facial regions. Despite this, MARE could still perceive, locate, and analyze the detected face image, thus generating more accurate and reliable reasoning content.